The first equation of each system is multiplied by 4. By what number would you multiply the second equation in order to eliminate the x variable by adding? First equation: 3x - 2y = 4 ; second equation: 3x - 2y = 4 ; second: 4x + 5y = 28
Answer by Fombitz(32388)

-3
.
.
.
12x-8y=16
-12x-15y=-84
Then adding,
12x-8y-12x-15y=16-84
-23y=-68